OpenStudy (anonymous):

determine the equation of the line that is tangent to y=2tanx, where x=pi/4

OpenStudy (anonymous):

\[y-y(x0)=m*(x-x0)\] where \[m=y'(x0)\] \[y'=(\sec(x))^{2} \] \[x0=pi/4\]

OpenStudy (anonymous):

sub pi/4 in the orignal equation to get y value?

OpenStudy (anonymous):

woah, can u show the steps to how u solved for x when sec^2x=0, and

OpenStudy (anonymous):

it's not zero sec²(pi/4)=2

OpenStudy (anonymous):